OpenStudy (anonymous):

Solve 2x2 = −10x. (4 points) x = 2, x = −5 x = 2, x = 5 x = 0, x = 5 x = 0, x = −5

OpenStudy (anonymous):

is these multiple choice?

OpenStudy (anonymous):

yep and its 2x^2

OpenStudy (anonymous):

ok it's gonna be x = 0 Or x = -5

OpenStudy (anonymous):

The last one

OpenStudy (anonymous):

k tnx

OpenStudy (anonymous):

2 x2 = -10 x 2 x2 +10 x = 0 2x (x+5) = 0 x = 0 OR x+5=0 x = 0 OR x=-5

OpenStudy (anonymous):

Does that help?